What is the mole fraction of Carbon Dioxide?

A. 0.0556

B. 0.0670

C. 0.2567

D. 0.6676

E. 0.7572


1
Expert's answer
2021-05-20T07:54:34-0400

Mole fraction of "CO_2" is

"8360 torr\times 1atm\over760 torr" "=11atm N_2"


"\\frac {101.1 kPa \u00d7 1 atm} {101.32} \n101.32\n101.1kPa\u00d71atm\n\u200b\t\n = 0.998 atm Ar"

"\\frac {760 mm Hg \u00d7 1 atm} {760 mm Hg} \n760mmHg\n760mmHg\u00d71atm\n\u200b\t\n = 1 atm CO2"

"Total Pressure = 11 atm + 5 atm + 0.998 atm +1 atm = 17.998 atm"

"mole fraction = \\frac {5 atm} {17.998} \n17.998\n5atm\n\u200b"

"Mole fraction of Carbon Dioxide= 0.2767" The correct answer is therefore "C"

.